Solve equations by factoring
Solve the equation.
x^2 + 2x - 8 = 0

Answer by ewatrrr(24785)


Hi,
Solving equation by factoring
x^2 + 2x - 8 = 0
(x+4)(x-2) = 0 Note:SUM of the inner product(4x) and the outer product(-2x) = 2x
(x+4)=0
x = -4
(x-2) = 0
x = 2